OpenStudy (anonymous):

PLEASE HELP?????!!!!!!!!!!!!!!!! Each year, Terrell buys a plane ticket to visit his grandmother for a week. Last year, he paid $132 . This year, he paid $162 . What was the percent increase in the cost of his plane ticket? Show your work.

OpenStudy (mertsj):

First find the increase. Now make a fraction. Put the increase on the top and the last year's amount on the bottom. Divide and change to a percent.

OpenStudy (anonymous):

WOULD THE FRACTION BE 30/100

OpenStudy (lgbasallote):

no...it would be \[\large \frac{\text{final value} - \text{inital value}}{\text{initial value}} \times 100\] therefore you get \[\large \frac{162 - 132}{132} \times 100\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

so... 30/132 times 100?